# WilKE: Wise-Layer Knowledge Editor for Lifelong Knowledge Editing

### v1

@@ Abstract
Knowledge editing aims to rectify inaccuracies in large language models (LLMs) without costly retraining for outdated or erroneous knowledge. However, current knowledge editing methods primarily focus on single editing, failing to meet the requirements for lifelong editing [^1] In this paper, lifelong editing is synonymous with lifelong knowledge editing.. This study reveals a performance degradation encountered by knowledge editing in lifelong editing, characterized by toxicity buildup and toxicity flash, with the primary cause identified as pattern unmatch. We introduce a knowledge editing approach named WilKE, which selects editing layer based on the pattern matching degree of editing knowledge across different layers. Experimental results demonstrate that, in lifelong editing, WilKE exhibits an average improvement of 46.2% and 67.8% on editing GPT2-XL and GPT-J relative to state-of-the-art knowledge editing methods.
@@ 1 Introduction
Large language models (LLMs) encode a wealth of world knowledge through pretraining on massive corpus (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023). However, outdated or erroneous knowledge may persist, and retraining these models with updated corpus incurs prohibitively high costs. To address this challenge, numerous studies have introduced knowledge editing (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022a; Meng et al., 2022b) as a solution, which involves updating the internal parameters of language models to edit specific knowledge.
@@
We analyze the primary reasons for these two phenomena, attributing them to pattern unmatch, as illustrated in Figure 2. Specifically, different layers of language model may detect different patterns, which is called key in key-value memories (Sukhbaatar et al., 2015; Sukhbaatar et al., 2019; Geva et al., 2020), thus extracting relevant information according to patterns and updating the hidden states. In other words, different knowledge may be stored in different layers, as illustrated in Section 4.3. However, ROME and MEMIT perform knowledge editing at predefined layers, which primarily lead to toxicity buildup and toxicity flash. To address this issue, we propose WilKE (Wise-Layer Knowledge Editor), which eliminates the need for predefined editing layer. Instead, WilKE selects editing layer based on the degree of pattern matching for different editing knowledge across various layers. Experimental results demonstrate that WilKE exhibits state-of-the-art comprehensive performance when editing GPT2-XL (1.5B) Radford et al. (2019) and GPT-J (6B) Wang and Komatsuzaki (2021). Specifically, in lifelong editing scenarios, under identical experimental conditions of conducting 1024 edits, WilKE demonstrates an average improvement of 46.2% and 67.8% in comprehensive performance relative to state-of-the-art methods when editing GPT2-XL and GPT-J, respectively.
@@
  • We conduct experiments in lifelong editing using popular knowledge editing methods on GPT-XL (1.5B) and GPT-J (6B), highlighting the superiority of WilKE over prevalent knowledge editing methods. The source code will be made available soon.
@@ 2 Related Work
Locate-and-edit methods first identify parameters corresponding to specific knowledge and achieve knowledge editing by updating these parameters. For example, Dai et al. (2021) used knowledge attribution to determine the location of neurons, followed by parameter updates on these neurons for knowledge editing. Meng et al. (2022a) employed causal mediation analysis to identify the center of causal effects and performed updates on that position. Meng et al. (2022b) extended upon ROME Meng et al. (2022a) by distributing residuals across multiple layers and achieved batch editing, and Li et al. (2023a) achieved more precise residual allocation.
However, existing knowledge editing methods are limited in single editing, overlooking their scalability. This oversight could lead to an overly optimistic view of knowledge editing methods and potential misuse when the methods are not yet mature. Therefore, for knowledge editing to be practically applicable, it is crucial to address the more challenging scenario, lifelong editing. Consequently, research into lifelong editing is imperative.
@@ 4.1 Toxicity
As editing progresses, the performance of the language model continuously deteriorates Yao et al. (2023), indicating that ongoing editing seems to introduce certain side effects. In this section, we refer to these side effects as "toxicity" and utilize rollback editing Li et al. (2023b) to define toxicity:
@@
Further investigation reveals that the data points causing toxicity flash exhibit the same "spike" phenomenon even in single editing. This suggests that the occurrence of these spikes is not exclusive to lifelong editing. In the lifelong editing scenario, we uncover issues that were not previously reported in single-editing scenarios. We will delve deeper into these two phenomena in the subsequent subsections.
@@ 4.3 Toxicity Flash
After further investigation, the fundamental reason lies in pattern unmatch, where the patterns of editing knowledge cannot be detected in the predefined editing layer. Continuing editing leads to language model overfitting, resulting in toxicity flash. The formal definition of pattern unmatch and experimental evidence are provided in Appendix D.
@@ 5.1 Where to Edit?
where $W_{fc}^{l}\in\mathbb{R}^{d\times d_{m}}$ and $W_{proj}^{l}\in\mathbb{R}^{d_{m}\times d}$ are the parameter matrices of the $l$-th layer’s FFN, $FFN^{l}$, and $d_{m}$ is the dimension of the intermediate hidden layer in the FFN. The symbol $\sigma$ represents the activation function, and $\boldsymbol{x}\in\mathbb{R}^{d}$ is the input to the FFN. As described in the key-value memories (Sukhbaatar et al., 2015; Sukhbaatar et al., 2019; Geva et al., 2020), $W_{fc}^{l}$ identifies patterns of the input $\boldsymbol{x}$ to obtain the key vector $k$, and then the value vector $v$ is retrieved from $W_{proj}^{l}$, as shown in Figure 2. Therefore, to achieve knowledge editing, we modify $W_{proj}^{l}$. After identifying the component that needs modification, we further determine the specific layer for modifying this component.
To find the editing layer $l^{*}$, our initial intuition is to identify the layer that produces the maximum activation strength for a specific knowledge, which is represented as $\text{argmax}_{l}\ \|\sigma(\boldsymbol{x}\cdot W_{fc}^{l})\|_{2}$. However, in reality, the optimization of $\boldsymbol{\delta}$ varies across different layers, as detailed in Appendix D.2. In other words, the importance of the hidden state outputted by different layers for editing specific knowledge is actually different. To comprehensively consider these two points, we define the editing layer (wise-layer) $l^{*}$ as follows:
$l^{*}=\text{argmin}_{l}\ \left\|\frac{\boldsymbol{\delta}}{\|W_{proj}^{l}\|_{2}\ \sigma(\boldsymbol{x}\cdot W_{fc}^{l})}\right\|_{2}$ (9)
@@ 5.2 How to Edit?
Same as Meng et al. (2022a), we introduce a residual term $\boldsymbol{\delta}\in\mathbb{R}^{d}$ to the output of the FFN in editing layer $l^{*}$, denoted as $FFN^{l^{*}}(\boldsymbol{x})+\boldsymbol{\delta}$. We optimize this residual term to align the model’s output with our expected results while not affecting irrelevant knowledge. For specific optimization details, we recommend interested readers to refer to Meng et al. (2022a).
Subsequently, we allocate the optimized residual $\boldsymbol{\delta^{*}}$ to $W_{proj}^{l}$ to accomplish knowledge editing:
$W_{proj}^{l^{*}}\leftarrow\frac{FFN(\boldsymbol{x})^{l^{*}}+\boldsymbol{\delta}^{*}}{\sigma(\boldsymbol{x}\cdot W_{fc}^{l^{*}})}$ (10)
@@ 6.2 Main Results
As shown in Table 1, we present the knowledge editing results after 1024 edits on GPT-XL and GPT-J. The results indicate that current knowledge editing methods perform poorly in lifelong editing, far from the optimistic results reported in single editing. However, these methods have been directly transferred and used in many other scenarios (Ma et al., 2023; Li et al., 2023a; Anonymous, 2024; Wang et al., 2023).
@@ 7 Conclusion
In this work, we focus on lifelong knowledge editing, finding that current knowledge editing methods suffer from severe performance degradation in lifelong editing. Our experimental results reveal the toxicity buildup and toxicity flash that may occur during lifelong editing, leading to the deterioration of model’s performance. The primary reason for these problems lies in pattern unmatch. To address this issue, we propose a model editing method called WilKE (Wise-Layer Knowledge Editor), which does not require predefined editing layer but selects editing layer based on specific editing knowledge. Experimental results demonstrate that in lifelong editing, WilKE achieves a significant improvement in overall performance compared to currently popular knowledge editing methods. In summary, our work is significant for improving knowledge editing methods and provide valuable insights for future work.
@@ 8 Limitation
Despite the promising performance of WilKE, our current studies still have limitations. Firstly, we select editing layer based on specific knowledge, yet knowledge may be distributed across multiple layers, leaving the question of how language models store knowledge is still under explored. Secondly, similar to previous knowledge editing research, we focus on factual knowledge assessment, which serves as a crucial entry point for our study on knowledge editing. Lastly, detecting match degree of specific knowledge across different layers of language models incurs a certain time cost, yet we believe this to be worthwhile in the initial stages of knowledge editing research.
@@ Appendix F Experimental Details
Reviewing Equation 6, here $\boldsymbol{y}_{e_{j}}^{\prime}$ may not be equal to $\boldsymbol{y}_{e_{j}}$, depending on whether the editing data after the test data will conflict with the existing knowledge Li et al. (2023b). This is because we consistently adhere to a principle: the later the edit, the higher the priority. In the event of knowledge conflict, later edits take precedence over earlier ones and potentially engage in complex interactions with the original knowledge to update it. For instance, as highlighted in Li et al. (2023b), if the model contains the fact "The notable work of Shakespeare is Hamlet" and undergoes the first edit "Hamlet was written in English normal-$\rightarrow$ French" followed by the second edit "Shakespeare wrote in French normal-$\rightarrow$ German" the second edit, if interacting with the original model’s fact, could result in a modification of the first edit’s outcome to "Hamlet was written in German" (though not modified explicitly in this way).
@@ F.1 Datasets
$Record^{E}$
The mother tongue of Danielle Darrieux is [French] $\Rightarrow$ [English].
$Record^{G}$
[Irrelevant Context]. Danielle Darrieux spoke the language [French] $\Rightarrow$ [English].
$Record^{L}$
The native language of Montesquieu is [French].
@@ F.2 Metrics
As previously mentioned, the issue of knowledge conflicts Li et al. (2023b) may arise in lifelong editing, potentially rendering the retention metric ineffective in the evaluation of lifelong editing methods Huang et al. (2023)Hartvigsen et al. (2022). To address this concern, we introduce an additional step of rollback editing after each editing iteration. Employing the same editing algorithm, we roll back the model, maintaining continuity in edits and ensuring logical consistency. Formally, after editing the model $f_{\theta_{i-1}}^{*}$ to obtain $f_{\theta_{i}}$, we denote the model after the rollback operation as $f_{\theta_{i}}^{*}$, and we expect the sequence $f_{\theta_{i}}^{*}\rightarrow f_{\theta_{i-1}}^{*}\rightarrow\cdots\rightarrow f_{\theta_{0}}^{*}$, where $f_{\theta_{0}}^{*}=f_{\theta_{0}}$.

### v2

@@ Abstract
Knowledge editing aims to rectify inaccuracies in large language models (LLMs) without costly retraining for outdated or erroneous knowledge. However, current knowledge editing methods primarily focus on single editing, failing to meet the requirements for lifelong editing [^1] In this paper, lifelong editing is synonymous with lifelong knowledge editing.. This study reveals a performance degradation encountered by knowledge editing in lifelong editing, characterized by toxicity buildup and toxicity flash, with the primary cause identified as pattern unmatch. We introduce a knowledge editing approach named Wise-Layer Knowledge Editor (WilKE), which selects editing layer based on the pattern matching degree of editing knowledge across different layers in language models. Experimental results demonstrate that, in lifelong editing, WilKE exhibits an average improvement of 46.2% and 67.8% on editing GPT2-XL and GPT-J relative to state-of-the-art knowledge editing methods.
@@ 1 Introduction
Large language models (LLMs) encode a wealth of world knowledge through pretraining on massive corpus (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023; Li et al., 2023a; Kale et al., 2023). However, outdated or erroneous knowledge may persist, and retraining these models with updated corpus incurs prohibitively high costs. To address this challenge, numerous studies have introduced knowledge editing (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022a; Meng et al., 2022b) as a solution, which involves updating the internal parameters of language models to edit specific knowledge.
@@
We analyze the primary reasons for these two phenomena, attributing them to pattern unmatch, as illustrated in Figure 2. Specifically, different layers of language model may detect different patterns, which is called key in key-value memories (Sukhbaatar et al., 2015; Sukhbaatar et al., 2019; Geva et al., 2020), thus extracting relevant information according to patterns and updating the hidden states. In other words, different knowledge may be stored in different layers, as illustrated in Section 4.3. However, ROME and MEMIT perform knowledge editing at predefined layers, which primarily lead to toxicity buildup and toxicity flash.
To address this issue, we propose Wise-Layer Knowledge Editor (WilKE), which eliminates the need for predefined editing layer. Instead, WilKE selects editing layer based on the degree of pattern matching for different editing knowledge across various layers. Experimental results demonstrate that WilKE exhibits state-of-the-art comprehensive performance when editing GPT2-XL (1.5B) Radford et al. (2019) and GPT-J (6B) Wang and Komatsuzaki (2021). Specifically, in lifelong editing scenarios, under identical experimental conditions of conducting 1024 edits, WilKE demonstrates an average improvement of 46.2% and 67.8% in comprehensive performance relative to state-of-the-art methods when editing GPT2-XL and GPT-J, respectively.
@@
  • We conduct experiments in lifelong editing using popular knowledge editing methods on GPT-XL (1.5B) and GPT-J (6B), highlighting the superiority of WilKE over prevalent knowledge editing methods. The source code is available at https://github.com/ChenhuiHu/WilKE.
@@ 2 Related Work
Locate-and-edit methods first identify parameters corresponding to specific knowledge and achieve knowledge editing by updating these parameters. For example, Dai et al. (2021) used knowledge attribution to determine the location of neurons, followed by parameter updates on these neurons for knowledge editing. Meng et al. (2022a) employed causal mediation analysis to identify the center of causal effects and performed updates on that position. Meng et al. (2022b) extended upon ROME Meng et al. (2022a) by distributing residuals across multiple layers and achieved batch editing, and Li et al. (2023b) achieved more precise residual allocation.
However, existing knowledge editing methods that modify model’s parameters mostly focus on single editing, unable to meet the demands of lifelong editing, leading to a certain gap between knowledge editing and practical applications. Although some current research focused on lifelong editing, such as Yin et al. (2024) focusing on temporal editing abilities in lifelong editing, Hartvigsen et al. (2022) and Huang et al. (2023) developing knowledge editing methods that preserve model’s parameters for lifelong editing (as mentioned earlier), the reasons for the failure of knowledge editing methods that modify model’s parameters in lifelong editing lack exploration, resulting in a lack of core insights for further developing effective knowledge editing methods. Consequently, research into lifelong editing is imperative.
@@ 4.1 Toxicity
As editing progresses, the performance of the language model continuously deteriorates Yao et al. (2023), indicating that ongoing editing seems to introduce certain side effects. In this section, we refer to these side effects as "toxicity" and utilize rollback editing Li et al. (2023c) to define toxicity:
@@
Additionally, toxicity flash is independent of editing order, which we explore in informal experiments, showing that even attempting to modify the order of editing, the phenomenon of toxicity flash persists at specific editing data points, accompanied by toxicity buildup. Further investigation reveals that the data points causing toxicity flash exhibit the same "spike" phenomenon even in single editing. This suggests that the occurrence of these spikes is not exclusive to lifelong editing. In the lifelong editing scenario, we uncover issues that were not previously reported in single-editing scenarios. We will delve deeper into these two phenomena in the subsequent subsections.
@@ 4.3 Toxicity Flash
4.4 Pattern Unmatch
After further investigation, the fundamental reason lies in pattern unmatch. Actually, pattern match and unmatch are relative concepts. As depicted in Figure 2, for the input "Thor is affiliated with the," there might not be any information extracted related to "Avengers" at layer $i+1$, but rather, the primary information about "Avengers" might be extracted at layer $i-2$. Therefore, layer $i-2$ effectively detects the pattern leading to the target output from the input, which is $\boldsymbol{k}^{i-2}$, representing as the key in key-value memory. Extracting the main information about "Avengers" , which is the target output, from $W_{proj}^{i-2}$ based on $\boldsymbol{k}^{i-2}$ and placing it in the residual flow constitutes what we refer to as pattern match. Otherwise, it is pattern unmatch. Thus, the ability to retrieve target information from $W_{proj}$ based on $\boldsymbol{k}$ serves as the criterion for determining whether the pattern matches.
In summary, the patterns of editing knowledge may not be detected in the predefined editing layer, which we call pattern unmatch. Continuing to edit on such knowledge will lead to language model overfitting, resulting in toxicity flash. The investigation details of pattern unmatch and experimental evidence are provided in Appendix D.
@@ 5.1 Where to Edit?
where $W_{fc}^{l}\in\mathbb{R}^{d\times d_{m}}$ and $W_{proj}^{l}\in\mathbb{R}^{d_{m}\times d}$ are the parameter matrices of the $l$-th layer’s FFN, $FFN^{l}$, and $d_{m}$ is the dimension of the intermediate hidden layer in the FFN. The symbol $\sigma$ represents the activation function, and $\boldsymbol{x}\in\mathbb{R}^{d}$ is the input to the FFN. As described in the key-value memories (Sukhbaatar et al., 2015; Sukhbaatar et al., 2019; Geva et al., 2020), $W_{fc}^{l}$ identifies patterns of the input $\boldsymbol{x}$ to obtain the key vector $\boldsymbol{k}^{l}$, and then the value vector $\boldsymbol{v}^{l}$ is retrieved from $W_{proj}^{l}$, as shown in Figure 2. Therefore, to achieve knowledge editing, we modify $W_{proj}^{l}$. After identifying the component that needs modification, we further determine the specific layer for modifying this component.
To find the editing layer $l^{*}$, our initial intuition is to identify the layer that produces the maximum activation strength for a specific knowledge, which is represented as $\text{argmax}_{l}\ \|\sigma(\boldsymbol{x}\cdot W_{fc}^{l})\|_{2}$. However, in practice, the optimization of $\boldsymbol{\delta}^{l}$ Meng et al. (2022a) after the FFN for aligning model’s output to achieve knowledge updating varies across different layers, as detailed in Appendix D.2. Specifically, $\boldsymbol{\delta}^{l}$ is calculated as follows:
$\begin{gathered}\boldsymbol{\delta}^{l}=\boldsymbol{v}^{l}_{*}-W_{proj}^{l}\boldsymbol{k}^{l},\boldsymbol{v}^{l}_{*}=\text{argmin}_{\boldsymbol{z}}\mathcal{L}(\boldsymbol{z})\\ \text{where }\mathcal{L}(\boldsymbol{z})=\frac{1}{N}\sum_{j=1}^{N}-\log\mathbb{P}_{f_{\theta}(\boldsymbol{m}_{i}^{l}:=\boldsymbol{z})}[o^{*}|c_{j}+p]\\ +D_{KL}(\mathbb{P}_{f_{\theta}(\boldsymbol{m}_{i}^{l}:=\boldsymbol{z})}[y|p^{\prime}]||\mathbb{P}_{f_{\theta}}[y|p^{\prime}])\end{gathered}$ (9)
where $\boldsymbol{m}_{i}^{l}$ represents the output of the MLP in the $l$th layer on the $i$th token, which is the end of the subject, $c_{j}$ denotes a token sequence randomly generated to simulate irrelevant context, $p$ corresponds to the knowledge we intend to edit, and $p^{\prime}$ represents the essence of the subject. In simple terms, the first term in $\mathcal{L}(\boldsymbol{z})$ is for knowledge updating, while the second term is for maintaining an understanding of the essence of the subject. Further details can be referenced in Meng et al. (2022a).
As mentioned before, the importance of the hidden state outputted by different layers for editing specific knowledge is actually different. To comprehensively consider these two points, namely the activation strength of specific knowledge and the $\boldsymbol{\delta}^{l}$ optimized for knowledge editing, we define the editing layer (wise-layer) $l^{*}$ as follows:
$l^{*}=\text{argmin}_{l}\ \left\|\frac{\boldsymbol{\delta}^{l}}{\|W_{proj}^{l}\|_{2}\ \sigma(\boldsymbol{x}\cdot W_{fc}^{l})}\right\|_{2}$ (10)
@@ 5.2 How to Edit?
Same as Equation 9, we introduce a residual term $\boldsymbol{\delta}^{l^{*}}\in\mathbb{R}^{d}$ to the output of the FFN in editing layer $l^{*}$, denoted as $FFN^{l^{*}}(\boldsymbol{x})+\boldsymbol{\delta}^{l^{*}}$. We optimize this residual term to align the model’s output with our expected results while not affecting irrelevant knowledge.
Subsequently, we allocate the optimized residual term $\boldsymbol{\delta}^{l*}$ to $W_{proj}^{l^{*}}$ to accomplish knowledge editing:
$W_{proj}^{l^{*}}\leftarrow\frac{FFN^{l^{*}}(\boldsymbol{x})+\boldsymbol{\delta}^{l^{*}}}{\sigma(\boldsymbol{x}\cdot W_{fc}^{l^{*}})}$ (11)
@@ 6.2 Main Results
As shown in Table 1, we present the knowledge editing results after 1024 edits on GPT-XL and GPT-J. The results indicate that current knowledge editing methods perform poorly in lifelong editing, far from the optimistic results reported in single editing. However, these methods have been directly transferred and used in many other scenarios (Ma et al., 2023; Li et al., 2023b; Anonymous, 2024; Wang et al., 2023).
@@ 7 Conclusion
In this work, we focus on lifelong knowledge editing, finding that current knowledge editing methods suffer from severe performance degradation in lifelong editing. Our experimental results reveal the toxicity buildup and toxicity flash that may occur during lifelong editing, leading to the deterioration of model’s performance. Through further investigation, we find that the direct cause of these problems lies in the predefined editing layer, while the underlying cause stems from pattern unmatch. To address this issue, we propose a model editing method called Wise-Layer Knowledge Editor (WilKE), which does not require predefined editing layer but selects editing layer based on the degree of pattern matching between different layers of the language model for specific editing knowledge. Experimental results demonstrate that in lifelong editing, WilKE significantly enhances overall performance compared to currently popular knowledge editing methods, achieving state-of-the-art knowledge editing performance. In summary, our work is significant for improving knowledge editing methods and provide valuable insights for future work.
@@ 8 Limitation
Despite the promising performance of WilKE, our current studies still have limitations. Firstly, we select editing layer based on specific knowledge, yet knowledge may be distributed across multiple layers, leaving the question of how language models store knowledge is still under explored. Secondly, similar to previous knowledge editing research, we focus on factual knowledge assessment, which serves as a crucial entry point for our study on knowledge editing. Furthermore, due to computational constraints, we did not conduct experiments on larger-scale language models but instead utilized GPT2-XL and GPT-J. However, WilKE does not require additional hypernetworks or other components, rendering it model-agnostic and thus exhibiting favorable scalability, enabling straightforward migration to larger models. Lastly, detecting match degree of specific knowledge across different layers of language models incurs a certain time cost, yet we believe this to be worthwhile in the initial stages of knowledge editing research.
@@ 9 Ethical Considerations
Acknowledgements
This work is supported by the Strategic Priority Research Program of Chinese Academy of Sciences (No. XDA27020203), the National Natural Science Foundation of China (No. 62176257).
@@ Appendix F Experimental Details
Reviewing Equation 6, here $\boldsymbol{y}_{e_{j}}^{\prime}$ may not be equal to $\boldsymbol{y}_{e_{j}}$, depending on whether the editing data after the test data will conflict with the existing knowledge (Li et al., 2023c; Jin et al., 2024a; Jin et al., 2024b). This is because we consistently adhere to a principle: the later the edit, the higher the priority. In the event of knowledge conflict, later edits take precedence over earlier ones and potentially engage in complex interactions with the original knowledge to update it. For instance, as highlighted in Li et al. (2023c), if the model contains the fact "The notable work of Shakespeare is Hamlet" and undergoes the first edit "Hamlet was written in English $\rightarrow$ French" followed by the second edit "Shakespeare wrote in French $\rightarrow$ German" the second edit, if interacting with the original model’s fact, could result in a modification of the first edit’s outcome to "Hamlet was written in German" (though not modified explicitly in this way).
@@ F.1 Datasets
$Record^{E}$ The mother tongue of Danielle Darrieux is [French] $\Rightarrow$ [English].
$Record^{G}$ [Irrelevant Context]. Danielle Darrieux spoke the language [French] $\Rightarrow$ [English].
$Record^{L}$ The native language of Montesquieu is [French].
@@ F.2 Metrics
As previously mentioned, the issue of knowledge conflicts Li et al. (2023c) may arise in lifelong editing, potentially rendering the retention metric ineffective in the evaluation of lifelong editing methods Huang et al. (2023)Hartvigsen et al. (2022). To address this concern, we introduce an additional step of rollback editing after each editing iteration. Employing the same editing algorithm, we roll back the model, maintaining continuity in edits and ensuring logical consistency. Formally, after editing the model $f_{\theta_{i-1}}^{*}$ to obtain $f_{\theta_{i}}$, we denote the model after the rollback operation as $f_{\theta_{i}}^{*}$, and we expect the sequence $f_{\theta_{i}}^{*}\rightarrow f_{\theta_{i-1}}^{*}\rightarrow\cdots\rightarrow f_{\theta_{0}}^{*}$, where $f_{\theta_{0}}^{*}=f_{\theta_{0}}$.